(Please show calculations for study purposes)

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 1.50 $6.20 $9.30

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 230 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 1,060 F
Labor spending variance $ 87 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week.
Answer #1

1. Labor efficiency variance = Labor spending variance - Labor rate variance

= -87-1060 = $1147 Unfavorable

Labor efficiency variance = (standard hours - Actual labor hours)*standard rate

-1147 = (230*1.50 - Actual labor hours)*6.20

Actual labor hours = 530 hours

2. Labor rate variance = (Standard rate - actual rate)*actual labor hours

1060 = (6.20 - actual hourly rate)*530

Actual hourly rate = $4.20 per hour